# Flow Matching for Tabular Data Synthesis

### v1

@@ 3.3 TabSynFlow: Flow Matching in Latent Spaces
After training, sampling proceeds by drawing $z_{0}\sim\mathcal{N}(0,I)$ and integrating the learned ODE until $t=1$. The resulting latent representation $z_{1}$ is then decoded to obtain a synthetic record $\tilde{x}$, followed by inverse transformation to the original feature space. Algorithms 12 in Appendix B.1 describe the procedure in detail.
@@ 4 Experimental Setting
Table 1 summarises the datasets used in our empirical evaluation. Following  ran2024multiobjective and  Little2024synthetic, we used four census datasets (from the UK, Fiji, Canada and Rwanda), plus one additional census data from Indonesia. We also added two further databases from UCI which are standard benchmarks for tabular data kotelnikov2023tabddpm. Unlike ran2024multiobjective, we treat age as continuous to exploit its natural ordering. Also, this choice remains compatible with TabDDPM/TabSyn, which operate on mixed-type data. To ensure a consistent and fair comparison, we adopt the same convention for TabSynFlow and TabbyFlow.
All the models in this study use a multilayer perceptron (MLP) architecture comprising four hidden layers with sizes [1024, 2048, 2048, 1024], configuration used by  zhang2024mixedtype. TabSyn and TabSynFlow employ Transformer-based architectures in their VAE counterparts, diverging from the standard MLP-based tabular VAE design (xu2019modeling).
Training was conducted with a batch size to 4096, and the ODE was integrated until $t=1$ using 100 steps  [^1] Unless otherwise specified, this is the default setup. Models were trained for up to 10,000 epochs, with early stopping based on the lowest observed training loss, which is similar to previous studies (zhang2024mixedtype; guzman-cordero2025exponential), although those studies capped training at 8000 epochs. The integration for synthesis uses Euler method. For sampling from the SDE, we set $g_{t}=\sigma_{t}$.
We focused the synthetic tabular data evaluation on preserving the statistical properties of the original data (utility) while minimising the potential to disclose sensitive information about individuals (disclosure risk) (taub2019the; elliot2023samples; Little2022comparing). Utility values range from 0 (no preservation of statistical properties) to 1 (perfect preservation), measuring how well synthetic data supports the same analyses as the original. Disclosure risk ranges from 0 (no added risk beyond baseline) to 1 (maximum disclosure risk), quantifying the probability that an adversary can correctly match sensitive information using synthetic data. Detailed evaluation methods are provided in Appendix C.
Table 1: Datasets used for the empirical study.
Abbr. Dataset Name #Obs. #Num. #Cat.
@@
UK UK Census 104267 1 14
CA Canada Census 32149 4 21
FI Fiji Census 84323 1 18
RW Rwanda Census 31455 1 12
ID Indonesia Census 177429 1 12
AD Adult 48842 5 10
CH Churn Modelling 10000 4 7
@@ 5 Results
In this section, we present and analyse the empirical findings of our study, highlighting the comparative performance of diffusion and flow-matching approaches for tabular data synthesis. As seen from the objective function, all algorithms optimise for utility; disclosure risk is measured post-hoc. We begin by establishing TabDDPM and TabSyn as strong baselines (Section 5.1). We then examine the number of function evaluations (NFEs), evaluating the utility and risk dynamics between computational efficiency (Section 5.2). Next, we investigate the impact of integration time ($t_{ode}$) on model performance, with specific attention to the evolution of probability paths (Section 5.3). Finally, we evaluate the effect of incorporating stochastic sampling within the VFM framework (Section 5.4). Collectively, these analyses provide a picture of how modelling choices shape the quality, efficiency, and reliability of synthetic tabular data generation.
@@ 5.1 Main Comparison
In this section, we present results for the census datasets which have been used for benchmarking in previous studies  (e.g. ran2024multiobjective; Little2024synthetic). Table 2 summarises the comparative performance of six synthetic data generation algorithms across UK, CA, FI, and RW datasets.
TabDDPM exhibits highly inconsistent results, excelling in UK while substantially underperforming in CA, FI, and RW. On the other hand, TabSyn demonstrates robust utility across all datasets, achieving values above $0.6$ in RW and exceeding $0.7$ in UK, CA, and FI, making it a reliable choice for utility-driven synthesis.
However, TabbyFlow proves to be the most effective model for these benchmarks. The OT trajectory delivers the highest utility in UK and CA, while the VP trajectory attains the best scores in FI and RW. Notably, TabbyFlow-VP consistently emerges as the second-best in utility, and further distinguishes itself by achieving lower risk, particularly in FI and RW.
From regular flow matching side, TabSynFlow-OT outperforms TabSyn in UK and RW, attaining lower disclosure risk in their high utility synthetic data, indicative of improved synthesis quality. In CA and FI, TabSynFlow demonstrates a classic risk–utility trade off. Moreover, TabSynFlow-VP achieves the most favorable risk profiles in CA, FI, and RW at the cost of lower utility.
In summary, TabbyFlow—with both OT and VP trajectories—consistently attains the highest overall average scores — calculated via average of utility and 1-risk — on the main census benchmark, achieving values of 0.7038 (OT) and 0.7031 (VP). TabSyn and TabSynFlow-OT also perform competitively, with average scores of 0.6742 and 0.6716, respectively, ranking just below the TabbyFlow models. In comparison, TabSynFlow-VP yields a lower average score of 0.6150, while TabDDPM lags behind at 0.5083. These findings underscore the robustness of flow matching algorithms, particularly the TabbyFlow family, in effectively high data utility and less disclosure risk. The findings also highlight the critical role of trajectory choice (OT or VP) that modulates the relationship in some datasets, particularly the risk as post-hoc consequence. Results for the additional datasets (Indonesia, Adult, and Churn) are available in Appendix D, and further analyses of alternative interpolants for both TabSynFlow and TabbyFlow are provided in Appendix E.2 and Appendix E.3, respectively.
Table 2: Comparative Performance of six synthetic data generation algorithms across four common census datasets (UK, CA, FI, RW). Bold and underline indicate the first and second best performing algorithms for each dataset, respectively. The table emphasises best performance of TabbyFlow in common census benchmarks.
@@ 5.2 Number of Function Evaluations between Tabsyn and Flow Matching Models
From the figure, it can be seen that the FM models generally converge after approximately 100 NFEs, indicating that using low steps in FM is sufficient. On the other hand, TabSyn requires more than 100 steps to outperform FM models in terms of utility, indicating a higher computational burden, in addition to costly training of VAE+diffusion.
In low-NFE settings, where flow-based models are designed to operate efficiently, TabSyn performed poorly. TabbyFlow-VP whilst yielding lower utility also achieves lower disclosure risk, and still outperforms TabSyn when NFEs are below 64. This early-stage advantage positions flow matching models, particularly using the OT path as promising candidates for applications operating under strict computational constraints.
In summary, while TabSyn can achieve competitive utility, it does so at the cost of significantly higher NFEs. Flow-matching models offer a more efficient alternative, especially in scenarios where computational efficiency is critical.
Figure 1: Average utility ($\uparrow$) and disclosure risk ($\downarrow$) as a function of the number of function evaluations (NFEs) for TabSyn and flow matching models (TabSynFlow-OT, TabSynFlow-VP, TabbyFlow-OT, TabbyFlow-VP) averaged across four datasets Solid lines represent utility, while dotted lines represent risk. Flow-matching models converge after approximately 100 NFEs, achieving competitive utility at substantially lower computational cost compared to TabSyn.
@@ 5.3 Flow Matching Results based on Integration Time
In contrast, both TabSynFlow and TabbyFlow using the VP path exhibit a lower initial utility at $t_{\text{ode}}=0.6$, suggesting that VP introduces disruptive noise early in the process. Utility and risk increase substantially when $t_{\text{ode}}>=0.6$, indicating that VP require full integration to achieve competitive utility. From these results, we can see that the consistent superiority of OT paths across integration times highlighting their robustness where early termination is desirable for computationally constrained applications. Dataset-specific results provided in Appendix D.4 show deviations around the average, particularly for $t_{\text{ode}}=[0.9,1]$, emphasising the importance of considering dataset characteristics in selecting integration parameters.
Figure 2: Average utility ($\uparrow$) and risk ($\downarrow$) against ODE integration time ($t_{\text{ode}}$) for TabSynFlow and TabbyFlow using OT and VP paths. Colours and line styles indicate method and path (VP light and solid, OT dark and dashed). OT paths enable early stopping without significant utility loss, whereas VP paths require full integration to achieve competitive performance.
@@ 5.4 SDE Implementation of Variational Flow Matching
In this subsection, we investigate whether stochastic differential equation (SDE) sampling yields results comparable to ODE sampling. Tables 4 and 4 present a comparative analysis of ODE and SDE formulations, with additional results provided in Appendix E.4 for alternative diffusion coefficients.
In general, it can be seen that the SDE approach performs as expected theoretically, producing marginal distributions similar to those generated by ODE  eijkelboom2024variationalflowmatchinggraph while, in some cases, producing superior performance. This is evidenced by the close alignment of utility and risk values across both formulations. This empirical validation confirms the mathematical foundation presented in the paper.
@@ 6.1 Discussion
Performance Comparison (Section 5.1). Flow matching models consistently outperform TabDDPM on most datasets, with TabbyFlow generating synthetic data of the highest utility in nearly all cases. This increased fidelity is generally accompanied by elevated disclosure risk, reflecting a positive empirical correlation between utility and risk. Our most interesting finding is while it is not commonly used in FM studies, using VP path in TabbyFlow has potential to produce better optimised synthetic data with higher utility but lower disclosure risk. TabSyn remains a reliable benchmark, while TabSynFlow-OT shows targeted improvements over TabSyn in selected cases but not universally. These results highlight flow matching—particularly TabbyFlow—as a viable alternative for tabular data synthesis, with trajectory choice gives different performance in terms of data utility and disclosure risk.
@@
Integration Time and Early Stopping (Section 5.3). OT paths outperform VP paths in both TabSynFlow and TabbyFlow, maintaining high utility earlier in the generation process. VP paths require full integration to achieve competitive performance. However, both trajectories are prone to reduced utility and / or increased risk when the integration time approaches one ($t_{ode}\to 1$). Practical strategies to mitigate this include early stopping ($t_{ode}\leq 0.9$) or velocity clipping222e.g., clip(1-1/(1-t), 0, m), m is the maximum value allowed. In general, OT trajectories provide a more stable and reliable approach to tabular synthesis.
@@ 6.2 Concluding Remarks
This study has presented a comprehensive evaluation of flow matching models for tabular data synthesis, comparing them with state-of-the-art diffusion-based approaches. Our results demonstrate that flow matching—particularly TabbyFlow—consistently achieves higher utility than diffusion baselines, while disclosure risk follows patterns determined by the empirical relationship between model fidelity and privacy exposure, rather than optimisation for risk reduction. Across diverse datasets, the OT trajectory consistently achieves balanced performance with high utility and reduced sensitivity to integration time, while the VP trajectory enhances privacy protection with lower risk at the expense of some utility. In practice, trajectory selection should be guided by specific application requirements.
TabbyFlow further distinguishes itself through its computational efficiency, achieving competitive results with less function evaluations (NFEs). In practice, we recommend the OT path as the default configuration, with early stopping ($0.9\leq t_{\text{ode}}<1.0$) and 64–128 NFEs to balance between synthetic data quality and computational cost. Even ultra-low settings (4-8 NFEs) remain viable for resource-constrained environments.
The introduction of stochasticity via SDE solvers produced nuanced outcomes. While SDE formulations theoretically preserve marginal distributions, empirical performance varied between datasets. In several cases, SDE sampling simultaneously improved utility and reduced risk, demonstrating that stochastic dynamics can act as a beneficial regulariser or enhance exploration of the data space. However, these benefits were not universal, indicating that the effectiveness of SDE is context-dependent. Therefore, SDE-based sampling should be considered an optional empirical extension, subject to dataset-specific validation.
Overall, flow matching with the OT paths emerges as an efficient approach to tabular data synthesis. Future work could explore hybrid trajectories, incorporate formal privacy guarantees such as differential privacy, and investigate more thoroughly the dataset-specific factors that drive the effectiveness of both ODE and SDE solvers.
Acknowledgments
B.I. Nasution is supported by the Indonesia Endowment Fund for Education Agency (LPDP) scholarship.

### v2

@@ 3.3 TabSynFlow: Flow Matching in Latent Spaces
After training, sampling proceeds by drawing $\tilde{z}_{0}\sim\mathcal{N}(0,I)$ and integrating the learned ODE until $t=1$. The resulting latent representation $\tilde{z}_{1}$ is then decoded to obtain a synthetic record $\tilde{x}$, followed by inverse transformation to the original feature space. Algorithms 12 in Appendix C.1 describe the procedure in detail.
@@ 4 Experimental Setting
4.1 Data
Table 1 summarises the datasets used in our empirical evaluation. Following  Ran et al. (2024) and  Little et al. (2024), we used four census datasets (from the UK, Fiji, Canada and Rwanda), plus one additional census data from Indonesia. We also added two further databases from UCI and Kaggle which are standard benchmarks for tabular data (Kotelnikov et al. 2023). Unlike Ran et al. (2024), we treat age as continuous to exploit its natural ordering. Also, this choice remains compatible with TabDDPM/TabSyn, which operate on mixed-type data. To ensure a consistent and fair comparison, we adopt the same convention for TabSynFlow and TabbyFlow. While these are our core benchmark datasets, we also provided stress test of TabSynFlow and TabbyFlow on high cardinality and dimensional data in Appendix G.2.
4.2 Model and Environments
To isolate the impact of the FM design axes studied in this paper (representation, learning target, trajectory, and dynamics), we control model capacity by using the same backbone across FM and diffusion baselines. Concretely, unless stated otherwise, all models are parameterised by a 4-layer MLP with hidden sizes $[1024,2048,2048,1024]$, following the mixed-type tabular synthesis setup of Zhang et al. (2024). This choice also aligns with the common practice in tabular diffusion, where the denoising network is typically a lightweight MLP (Kotelnikov et al. 2023; Zhang et al. 2024).
For latent-space methods, the encoder and decoder operate on the original mixed-type table and are kept as in the corresponding reference implementations. In particular, TabSyn and TabSynFlow employ Transformer-based VAE components for tokenisation and de-tokenisation, while the diffusion or FM model is trained in the latent space using the same MLP setting. For TabbyFlow, prior work uses transformer + MLP backbones (Guzmán-Cordero et al. 2025). In our benchmark, we additionally report an architecture ablation for TabbyFlow (MLP vs Transformer) to quantify sensitivity to the backbone choice while preserving comparability with MLP-based baselines (Appendix G.3).
Training was conducted with a batch size to 4096, and the ODE was integrated until $t=1$ using 100 steps  [^1] Unless otherwise specified, this is the default setup. Models were trained for up to 10,000 epochs, with early stopping based on the lowest observed training loss, which is similar to previous studies (Zhang et al. 2024; Guzmán-Cordero et al. 2025), although those studies capped training at 8000 epochs. The integration for synthesis uses Euler method. For sampling from the SDE, we set $g_{t}=\sigma_{t}$. All reported numbers are averaged over 20 random seeds of data synthesis. Error bars in tables denote $\pm 1$ standard deviation across seeds. For computational resource, we run all experiments using Python 3.11 with library torch 2.3.0 on an L40S GPU.
Table 1: Datasets used for the study. $\sum K_{d}$ is the sum of category counts over categorical variables, while $\max(K_{d})$ is the highest cardinality among categorical variables. These statistics summarise both overall categorical complexity and the hardest single-variable regime.
Abbr. Dataset Name #Obs. #Numerical #Categorical $\sum K_{d}$ $\max(K_{d})$
@@ 4.2 Model and Environments
UK UK Census 104267 1 14 218 74
CA Canada Census 32149 4 21 170 37
FI Fiji Census 84323 1 18 239 37
RW Rwanda Census 31455 1 12 197 37
ID Indonesia Census 177429 1 12 62 10
AD Adult 48842 5 10 122 42
CH Churn Modelling 10000 4 7 26 11
4.3 Evaluation Metrics
We evaluate synthetic microdata using two criteria widely adopted in official statistics and statistical disclosure control: Utility (Section 4.3.1) and Disclosure Risk (Section 4.3.2(Taub et al. 2020; Little et al. 2022). Our evaluation is motivated by the common practice of tabular microdata, especially in NSOs, for decision-making through interpretable analytical outputs. To support our utility and risk analysis, we also provided additional metrics on data fidelity (Section 4.3.3). Full metric definitions and additional results are provided in Appendix D.
4.3.1 Utility
Utility refers to how well synthetic data supports the same practical analyses as the real data, such as producing descriptive statistics and fitting statistical models. In this paper, we evaluate utility through two common uses in official statistics: tabulation and inference. Tabulation utility is measured using the ratio-of-counts (ROC) score on univariate and bivariate frequency tables. Inferential utility is measured using confidence interval overlap (CIO) for regression coefficients, a standard task-specific utility metric for synthetic data (Ran et al. 2024; Elliot et al. 2023; Little et al. 2021). We aggregate these three components as:
$\text{Utility}=\frac{\mathrm{ROC_{uni}}+\mathrm{ROC_{biv}}+\mathrm{CIO}}{3}.$ (4.1)
4.3.2 Disclosure Risk
We measure attribute disclosure risk using the targeted correct attribution probability (TCAP) framework for synthetic microdata (Taub et al. 2019; Little et al. 2022). TCAP considers an intruder who (i) knows the individual appears in the original dataset, (ii) knows that individual’s values on a set of quasi-identifying key variables, and (iii) searches the synthetic data for records matching those keys to infer a sensitive target attribute (Little et al. 2024). For each synthetic record $j$ with key and target values $(K^{\prime}_{j},T^{\prime}_{j})$, we compute the empirical conditional probability in the original data (see Appendix D.2), and aggregate these values across synthetic records. Smaller TCAP indicates a lower probability of correctly inferring the target attribute and therefore lower attribute disclosure risk.
4.3.3 Additional Metrics
In addition to these primary scores, we also compute widely used diagnostic metrics in tabular deep generative models for data fidelity (Zhang et al. 2024; Guzmán-Cordero et al. 2025; Mueller et al. 2025). Data fidelity is defined as the extent to which the synthetic data is statistically indistinguishable from the real data, both at the level of marginal distributions and multivariate structure Adams et al. (2025). This study calculates data fidelity using column-wise Wasserstein distance for continuous attributes, total variation distance for categorical attributes, low-order shape and trend statistics, a classifier two-sample test (C2ST) detection score, and sample-level $\alpha$-precision and $\beta$-recall.
@@ 5 Results
In this section, we present and analyse the empirical findings of our study, highlighting the comparative performance of diffusion and flow-matching approaches for tabular data synthesis. As seen from the objective function, all algorithms optimise for utility; disclosure risk is measured post-hoc. We begin by comparing investigating FM methods performance on tabular data synthesis, especially compared to diffusion baselines (Section 5.1). We then examine the number of function evaluations (NFEs), evaluating the utility and risk dynamics between computational efficiency (Section 5.2). Next, we investigate the impact of integration time ($t_{\text{ode}}$) on model performance, with specific attention to the evolution of probability paths (Section 5.3). Finally, we evaluate the effect of incorporating stochastic sampling within the VFM framework (Section 5.4). Collectively, these analyses provide a picture of how modelling choices shape the quality, efficiency, and reliability of synthetic tabular data generation.
@@ 5.1 Main Comparison
We report results across seven datasets (five census-type and two standard tabular benchmarks) comparing diffusion baselines (TabDDPM, TabSyn) against flow matching variants in latent space (TabSynFlow-OT/VP) and data space (TabbyFlow-OT/VP). All metrics are reported as mean $\pm$ standard deviation across 20 seeds or data synthesis.
Across the four census benchmarks (UK, CA, FI, RW), flow matching is consistently competitive with diffusion, and TabbyFlow achieves the strongest utility overall. Trajectory choice shifts the operating point: TabbyFlow-OT performs best on UK and CA, whereas TabbyFlow-VP performs best on FI and RW, indicating that VP can improve utility in some census settings while maintaining low risk. TabSyn remains a strong baseline, and TabSynFlow often matches or improves TabSyn under OT, with VP typically reducing risk at the cost of utility.
The same pattern largely carries over to the three additional datasets (ID, Adult, Churn). TabbyFlow-OT achieves the highest utility on ID and Adult, while TabSynFlow-OT is strongest on Churn, suggesting that latent-space FM can be preferable when the dataset structure favours compact representation learning. Risk is low across most settings but remains dataset-dependent: VP generally yields lower risk than OT for a comparable method, consistent with its more conservative sampling behavior. Overall, the results support two practical conclusions: (i) flow matching provides performance comparable to or better than diffusion for mixed-type tabular synthesis, and (ii) the choice between latent-space and data-space FM, as well as OT versus VP trajectories, offers a controllable utility–risk trade-off that varies with dataset characteristics.
Table 2: Comparative Performance of six synthetic data generation algorithms across four common census datasets (UK, CA, FI, RW). We report the mean $\pm$ standard deviation across 20 seeds of synthetic data generation. Bold and underline indicate the first and second best performing algorithms for each dataset, respectively. The table emphasises best performance of TabbyFlow in common census benchmarks.
@@
[…]
Table 3: Comparative Performance of six synthetic data generation algorithms across three datasets (ID, AD, CH). We report the mean $\pm$ standard deviation across 20 seeds of synthetic data generation. Bold and underline indicate the first and second best performing algorithms for each dataset, respectively. The table highlights the TabbyFlow and TabSynFlow-OT which stands out as a strong alternative to TabSyn.
Evaluation Algorithm ID AD CH
@@ 5.2 Number of Function Evaluations between Tabsyn and Flow Matching Models
From the figure, it can be seen that the FM models generally converge after approximately 100 NFEs, indicating that using low steps in FM is sufficient. On the other hand, TabbyFlow outperforms TabSyn in all NFE settings based on the utility. Furthermore, TabSyn requires more than 100 steps to outperform FM models, especially TabSynFlow-OT in terms of utility, indicating a higher computational burden, in addition to costly training of VAE+diffusion.
In low-NFE settings, where flow-based models are designed to operate efficiently, TabSyn performed poorly. TabbyFlow-VP whilst yielding lower utility also achieves lower disclosure risk, and still outperforms TabSyn when NFEs are $\leq$ 32. Meanwhile, TabSynFlow has synthetic data with lowest disclosure risk, while also having higher utility than TabSyn when NFEs are $\leq$ 16. This early-stage advantage positions flow matching models as promising candidates for applications operating under strict computational constraints. In summary, while TabSyn can achieve competitive utility, it does so at the cost of significantly higher NFEs. Flow-matching models offer a more efficient alternative, especially in scenarios where computational efficiency is critical.
((a))
((b))
Figure 1: Average utility ($\uparrow$) and disclosure risk ($\downarrow$) as a function of the number of function evaluations (NFEs). Results are averaged across four datasets. (a) Overall trends shown on the full scale $[0,1]$. (b) Zoomed view of the disclosure risk to highlight differences between methods. TabSyn (blue, solid) is compared against flow-matching models (TabSynFlow-OT, TabSynFlow-VP, TabbyFlow-OT, TabbyFlow-VP). For flow-matching models, colours and line styles indicate the probability path (VP: lighter colours and solid lines; OT: darker colours and dashed lines). The vertical dashed line marks 100 NFEs, corresponding to the default implementation. Across datasets, flow-matching models saturate after approximately 100 NFEs, achieving competitive utility while requiring substantially fewer function evaluations than TabSyn.
@@ 5.3 Flow Matching Results based on Integration Time
In contrast, both TabSynFlow and TabbyFlow using the VP path (solid and lighter lines) exhibit a lower initial utility at $t_{\text{ode}}=0.6$, suggesting that VP introduces disruptive noise early in the process. Utility and risk increase substantially when $t_{\text{ode}}>=0.6$, indicating that VP require full integration to achieve competitive utility. From these results, we can see that the consistent superiority of OT paths across integration times highlighting their robustness where early termination is desirable for computationally constrained applications. Dataset-specific results provided in Appendix E.3 show deviations around the average, particularly for $t_{\text{ode}}=[0.9,1]$, emphasising the importance of considering dataset characteristics in selecting integration parameters.
Figure 2: Average utility ($\uparrow$) and risk ($\downarrow$) against ODE integration time ($t_{\text{ode}}$) for TabSynFlow (green) and TabbyFlow (red) using OT and VP paths. Colours and line styles indicate method and path (VP light and solid, OT dark and dashed). OT paths enable early stopping without significant utility loss, whereas VP paths require full integration to achieve competitive performance.
@@ 5.4 SDE Implementation of Variational Flow Matching
In this subsection, we investigate whether stochastic differential equation (SDE) sampling yields results comparable to ODE sampling. Tables 5 and 5 present a comparative analysis of ODE and SDE formulations, with additional results provided in Appendix F.5 for alternative diffusion coefficients.
In general, it can be seen that the SDE approach performs as expected theoretically, producing marginal distributions similar to those generated by ODE  Eijkelboom et al. (2024) while, in some cases, producing superior performance. This is evidenced by the close alignment of utility and risk values across both formulations. This empirical validation confirms the mathematical foundation presented in the paper.
@@ 6.1 Discussion
Performance Comparison (Section 5.1). Flow matching models consistently outperform TabDDPM on most datasets, with TabbyFlow generating synthetic data of the highest utility in nearly all cases. This increased utility is generally accompanied by elevated disclosure risk, reflecting a positive empirical correlation between utility and risk. An interesting finding is that using a VP path in TabbyFlow can yield higher utility with lower disclosure risk, even though VP paths are not commonly used in FM studies. TabSyn remains a reliable benchmark, while TabSynFlow-OT shows targeted improvements over TabSyn in selected cases but not universally. These results highlight flow matching—particularly TabbyFlow—as a viable alternative for tabular data synthesis, with trajectory choice gives different performance in terms of data utility and disclosure risk.
@@
Integration Time and Early Stopping (Section 5.3). OT paths outperform VP paths in both TabSynFlow and TabbyFlow, maintaining high utility earlier in the generation process. VP paths require full integration to achieve competitive performance. However, both trajectories are prone to reduced utility and / or increased risk when the integration time approaches one ($t_{\text{ode}}\to 1$). Practical strategies to mitigate this include early stopping ($t_{\text{ode}}\leq 0.9$) or velocity clipping222e.g., clip(1-1/(1-t), 0, m), m is the maximum value allowed. In general, OT trajectories provide a more stable and reliable approach to tabular synthesis.
@@ 6.2 Concluding Remarks
This study has presented a comprehensive evaluation of flow matching models for tabular data synthesis, comparing them with state-of-the-art diffusion-based approaches. Flow matching, particularly TabbyFlow, achieves higher utility across datasets, while disclosure risk follows patterns determined by the empirical relationship between model utility and privacy, rather than optimisation for risk reduction. OT trajectories provide a strong default with stable utility and low sensitivity to integration time, whereas VP typically lowers risk at some utility cost, so the choice should depend on application requirements.
TabbyFlow is also computationally efficient, reaching competitive performance with fewer function evaluations. We recommend OT with early stopping ($0.9\leq t_{\text{ode}}<1.0$) and 64–128 NFEs as a practical quality–cost trade-off, with 4–8 NFEs remaining usable in resource-constrained settings. SDE sampling can sometimes improve both utility and risk, but its benefits are dataset-dependent, so it is best treated as an optional extension validated per dataset.
Overall, both latent-space and data-space flow matching (including variational formulations) are effective for tabular synthesis in our study. Future work should expand evaluation beyond census-style tables to higher-dimensional domains (for example, health or genomics), study mixed continuous–categorical generation more directly (including hybrid conditional and discrete FM (Gat et al. 2024)), and incorporate formal privacy guarantees such as differential privacy for sensitive deployments.
Broader Impact Statement
Synthetic tabular microdata can help statistical agencies and other data custodians broaden access to data for research and policy analysis when releasing unit records is restricted. However, disclosure risk can persist through linkage or attribute inference, and low-fidelity synthesis can bias results, especially for small subpopulations or rare categories. This work provides empirical evidence on how flow-based design and sampling choices shift the utility–risk relation under evaluation practices used in official statistics. Any deployment should be dataset-specific and include domain review, clear communication of intended use, and risk assessment reported alongside utility. These considerations also apply to other sensitive tabular domains such as health, finance, and public services.